# Multiply Robust Estimation for Local Distribution Shifts with Multiple Domains

### v1

@@ 6 Experiments
We evaluate the methods on both simulated and real data. We use mean squared error (MSE) for regression tasks and cross entropy (CE) for classification tasks for evaluation. Standard errors are reported in parentheses in the following tables. More implementation details such as weights estimation and model hyperparameters are included in Section C.2. The code is available at github.com/MREstimation/MR.
@@
In summary, we find that the proposed MR consistently outperforms the competitors. In particular, MR achieves the lowest test losses on the majority of segments (see results in Figure 3 and Appendix C) on all datasets. Since XGB and DR does not account for the distributional heterogeneity across segments, their performances are typically sub-optimal. The improvement of DR-SF over DR in most cases suggests that including one-hot segment feature helps the model adapt to different segments, though with limited capability. The performance of MFSAN and DADIL are sensitive to hyperparameter settings. Even after fine-tuning (see Section C.3 for more details), they still cannot achieve comparable performance to MR and DR-SF. The issue with MFSAN lies in its attempt to align the predictions made by all the classifiers learned from different sources (Equation (6) in their paper). This alignment penalty in MFSAN could potentially hinder its performance, as it contradicts the goal of leveraging the heterogeneity of data from various sources. On the other hand, the dictionary learning approach, which is the core of DADIL, assumes that the data distribution in each target domain can be well represented by a mixture of learned distributions. However, this assumption is restrictive and may not hold in real-world applications. Even if this assumption holds, effectively learning the dictionary of moderate to high dimensional densities becomes challenging due to the curse of dimensionality. In addition, it is worth noting that MFSAN and DADIL require significantly more computation time compared to other competing methods even with the use of GPU [^4] Details on computing environment are in Section C.3 (Table 2).
@@ 6.2 Public Datasets
6.3 User City Prediction
We consider a dataset utilized for user city prediction from a large technology company. The observational unit of the dataset is a (user, city) pair and the features are mostly based on the interaction between a user and a city. A user could have multiple candidate cities and thus be associated with multiple (user, city) pairs. The labels are collected from the users, based on responses to a survey. The responses $y\in\{0,1\}$ indicate whether a given city is the primary location of a given user. These predictions are then fed through downstream applications for multiple use cases.
@@ 6.3 User City Prediction
We assess each method’s performance using test CE relative to the production model. We also include the Brier score, a measure of calibration of a classifier (Brier, 1950). As shown in Table 2, our multiply-robust method consistently outperforms others in CE and Brier score. The CE results by country are also visualized in Figure 3, demonstrating improved performance compared to the production model across nearly all segments.
@@ 7 Conclusion and Future Work
Existing assumptions for domain adaptation can be overly restrictive when applied to an entire population. To address this, we introduce and study a domain adaptation framework that allows for a relaxed assumption of classical distribution shifts only restricted to segments of the data. This presents a statistical challenge in refining a predictor across all segments of our population.
To address this challenge, we present a method for refining models across segments within a population. We provide both theoretical justification and empirical evidence demonstrating the enhanced performance of our method compared to existing state-of-the-art domain adaptation techniques. Our approach is easily implementable using off-the-shelf algorithms.
Potential avenues for future work include exploring methods of jointly learning each of the refined models, by incorporating segmentation clustering, treating this as a multitask learning problem. Extensions to image and text data could also be considered, possibly utilizing pre-trained models in place of the base models in Algorithm 1. These problems however would involve further consideration of the trade-off between domain invariance properties and refinement for each segment.
@@ A.1 Methods for Learning Importance Weights
We can apply standard techniques for learning the importance weights of the function class. Firstly, we consider learning covariate shift weights through discriminative learning, i.e. learning a probabilistic classifier that discerns between the weights (Bickel et al., 2009). Let $T=\{0,1\}$ denote whether a sample is within the training or test distribution letting $P(\cdot)=\mathbb{P}(\cdot|T=0)$ and $Q(\cdot)=\mathbb{P}(\cdot|T=1)$, then the importance weights can be expressed up to proportion
$w(x)=\frac{dQ(x)}{dP(x)}\propto\frac{\mathbb{P}(T=1|x)}{\mathbb{P}(T=0|x)}$
@@
Secondly, we consider Penalized Risk Minimization (Nguyen et al., 2010) Penalized Risk Minimization involves the estimation as follows:
$\widehat{w}(x)$ $=\widehat{g}(x)$
$\widehat{g}$ $=\text{argmin}_{g\in\mathcal{G}}\frac{1}{n}\sum_{i=1}^{n}g(X_{i})$
$-\frac{1}{n}\sum_{i=1}^{n}\log(g(X^{\prime}_{i}))+\frac{\gamma_{n}}{2}I^{2}(g)$
where $\mathcal{G}$ is a reproducing Kernel Hilbert space and $I$ is a non-negative measure of complexity for $g$ such that $I(w)<\infty$ (for example, $I(g)=\lVert g\rVert_{\mathcal{F}}$).
Another common approach to estimating importance weights is via Kernel Mean Matching (KMM). Given a positive definite kernel function $k(\cdot,\cdot)\to\mathbb{R}$ let $G_{ij}=k(X_{i},X_{j})$ and let $\kappa_{i}=\frac{n_{tr}}{n_{te}}\sum_{j=1}^{n_{te}}k(X_{i},X_{j}^{\prime})$. Then kernel mean matching computes a set of weights $\{w_{i}\}_{i=1}^{n_{tr}}$
@@
Lastly, when correcting for label shift one can use the Black Box Shift Estimation (BBSE) of (Lipton et al., 2018). This involves taking a predictor $h(\cdot)$, such that $\widehat{Y}=h(X)$. The corresponding importance weights $\mathbf{\tilde{w}}\in\mathbb{R}^{K}$ where $\tilde{w}(y)=\frac{Q(y)}{P(y)}$, can be defined through a linear equation of the expected confusion matrix $\mathbf{C}$, where $\mathbf{C}_{ij}=P(h(X_{i}):=\widehat{Y}=i,Y=j)$ and the predicted class frequencies on the test set $\mathbf{\mu}_{Q,h}\in\mathbb{R}^{K}$ where $\mu_{Q,h,j}=Q(\widehat{Y}=j)$. We differentiate the importance weights for each class $\mathbf{\tilde{w}}\in\mathbb{R}^{K}$ from $\mathbf{w}\in\mathbb{R}^{n}$, the weight vector for each training observation. The importance weights can then be defined by the solution to the following:
@@
Using the training labels, one can estimate $\mathbf{\widehat{\mu}}_{Q,h}$ and $\mathbf{\widehat{C}}$ using empirical means and as long as $\sigma_{min}$, the smallest eigenvalue of $\mathbf{C}$ is invertible, then one can estimate the importance weights of the class
@@
B.1 Proof of Theorem 5.3
@@
We consider the two-stage multiply robust estimator
$\widehat{f}_{MR}^{(s)}$ $=\text{argmin}_{f\in\mathcal{F}}R[f|Y,X,s;\widehat{w}_{s}]+\lambda^{\prime}\lVert\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}-f\rVert_{\mathcal{F}}$
$=\text{argmin}_{f:\lVert\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}-f\rVert_{\mathcal{F}}\leq\nu^{\prime}}R[f|Y,X,s;\widehat{w}_{s}].$
The constraint value $\nu^{\prime}$ denotes the corresponding value to the regularization parameter $\lambda^{\prime}$.
@@
$\sup_{f:\lVert f-\mathbf{h}^{\intercal}\widehat{\beta}\rVert\leq\nu^{\prime}}\left|R[f|Y,X,s;w_{s}]-R_{Q}[f|s]\right|$ $\leq B(n_{s},\delta)$
$=\eta L\left(\frac{2\nu^{\prime}}{n_{s}}\sqrt{\text{tr}(\mathbf{K}_{s})}+3\sqrt{\frac{\log(2/\delta)}{n_{s}}}+\frac{1}{n_{s}}\sqrt{\sum_{i\in N_{s}}(\mathbf{h}^{\intercal}\widehat{\beta})^{2}(X_{i})}\right)$
@@
$|R_{Q}[f|Y,X,s,\widehat{w}_{s}]-R_{Q}[f|Y,X,s,w_{s}]|\leq C(n_{s},\widehat{w}_{s})=L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
@@
$R_{Q}[\widehat{f}_{MR}^{(s)}|s]$ $\leq R_{Q}[\widehat{f}_{MR}^{(s)}|Y,X,s,w_{s}]+B(n_{s},\delta)$
$\leq R[\widehat{f}_{MR}^{(s)}|Y,X,s,\widehat{w}_{s}]+B(n_{s},\delta)+C(n_{s},\widehat{w}_{s})$
$\leq R[f^{*(s)}_{\nu_{q,s}}|Y,X,s,\widehat{w}_{s}]+B(n_{s},\delta)+C(n_{s},\widehat{w}_{s})$
$\leq R[f^{*(s)}_{\nu_{q,s}}|Y,X,s,w_{s}]+B(n_{s},\delta)+2C(n_{s},\widehat{w}_{s})$
@@
The third inequality follows from the fact $R[\widehat{f}_{MR}^{(s)}|Y,X,s,\widehat{w}_{s}]\leq R[f^{*(s)}_{\nu_{q,s}}|Y,X,s,\widehat{w}_{s}]$
@@
B.2 Supporting Lemmas
Lemma B.1 (Best Linear Combination As A Prior).
[…]
For a universal constant $C$
Lemma B.1 follows from a simple application of the results from (Ostrovskii & Bach, 2021). The main idea we exploit is that learning a simple linear combination is easier than learning in the larger function class $\mathcal{F}$. Since we assume that the total number of base estimators is not increasing, then $M_{eff}$, $M_{r}$ and $\tilde{\lambda}_{1}$ are all constants and we recover the stage 1 model at a rate of $n_{s}^{-1/2}$. This therefore, allows us to learn this prior faster than that of the stage 2 model.
B.3 Proof of Lemma B.1
Proof.
Let $h_{m}:\mathcal{X}\to\mathcal{Y}$ denote a set of base-model functions, and let $\mathbf{h}$ denote the vector of such functions, such that $\mathbf{h}^{\intercal}\beta=\sum_{m=1}^{M}h_{m}\beta_{m}$.
The first step is an immediate application of Theorem 1.1 of (Ostrovskii & Bach, 2021) letting $\rho$ denote the minimum value such that $\Sigma\preceq\rho H$. Then for all $n_{s}\geq\rho M_{eff}M\log(eM/\delta)$
$(\widehat{\beta}^{(s)}-\beta^{*(s)})^{\intercal}H^{2}(\widehat{\beta}^{(s)}-\beta^{*(s)})\lesssim\frac{M_{eff}\log(e/\delta)}{n_{s}}$
where $\lesssim$ denotes $\leq$ up to a universal constant $C$.
Further let $\tilde{\lambda}_{1}$ denote the smallest eigenvalue of $H$. Then with probability at least $1-\delta$.
$\lVert\widehat{\beta}^{(s)}-\beta^{*(s)}\rVert_{2}^{2}\lesssim\frac{M_{eff}\log(e/\delta)}{\lambda^{2}_{1}n_{s}}$
$\lVert\mathbf{h}^{\intercal}\beta^{*(s)}-\mathbf{h}^{\intercal}\widehat{\beta}\rVert_{\mathcal{F}}$ $\leq\sum_{m=1}^{M}\lVert h_{m}(\beta_{m}^{*(s)}-\widehat{\beta}_{r})\rVert_{\mathcal{F}}$

### v2

@@ 6 Experiments
We evaluate the methods on both simulated and real data. We use mean squared error (MSE) for regression tasks and cross entropy (CE) for classification tasks for evaluation. Standard errors are reported in parentheses in the following tables. More implementation details such as weights estimation and model hyperparameters are included in Section C.2. The code is available at github.com/facebookresearch/multiply_robust.
@@
In summary, we find that the proposed MR consistently outperforms the competitors. In particular, MR achieves the lowest test losses on the majority of segments (see results in Figure 3 and Appendix C) on all datasets. Since XGB and DR do not account for the distributional heterogeneity across segments, their performances are typically sub-optimal. The improvement of DR-SF over DR in most cases suggests that including one-hot segment feature helps the model adapt to different segments, though with limited capability. The performance of MFSAN and DADIL are sensitive to hyperparameter settings. Even after fine-tuning (see Section C.3 for more details), they still cannot achieve comparable performance to MR and DR-SF. The issue with MFSAN lies in its attempt to align the predictions made by all the classifiers learned from different sources (Equation (6) in their paper). This alignment penalty in MFSAN could potentially hinder its performance, as it contradicts the goal of leveraging the heterogeneity of data from various sources. On the other hand, the dictionary learning approach, which is the core of DADIL, assumes that the data distribution in each target domain can be well represented by a mixture of learned distributions. However, this assumption is restrictive and may not hold in real-world applications. Even if this assumption holds, effectively learning the dictionary of moderate to high dimensional densities becomes challenging due to the curse of dimensionality. In addition, it is worth noting that MFSAN and DADIL require significantly more computation time compared to other competing methods even with the use of GPU [^4] Details on computing environment are in Section C.3 (Table 2).
@@ 6.2 Public Datasets
6.3 Meta User City Prediction
We consider a dataset utilized for user city prediction from Meta. The observational unit of the dataset is a (user, city) pair and the features are mostly based on the interaction between a user and a city. A user could have multiple candidate cities and thus be associated with multiple (user, city) pairs. The labels are collected from the users, based on responses to a survey. The responses $y\in\{0,1\}$ indicate whether a given city is the primary location of a given user. These predictions are then fed through downstream applications for multiple use cases.
@@ 6.3 Meta User City Prediction
We assess each method’s performance using test CE relative to the production model. We also include the Brier score, a measure of calibration of a classifier (Brier, 1950). As shown in Table 2, our multiply robust method consistently outperforms others in CE and Brier score. The CE results by country are also visualized in Figure 3, demonstrating improved performance compared to the production model across nearly all segments.
@@ 7 Conclusion and Future Work
Existing assumptions for domain adaptation can be overly restrictive when applied to an entire population. To address this, we introduce and study a domain adaptation framework that allows for the relaxation of classical distribution shift assumptions to hold only within segments of the data. This presents a statistical challenge in refining a predictor across all segments of our population.
To address this challenge, we present a method to refine models across population segments. We provide theoretical justification and empirical evidence demonstrating the enhanced performance of our method compared to existing state-of-the-art domain adaptation techniques. Our approach is easily implementable using off-the-shelf algorithms.
Potential avenues for future work include exploring methods of jointly learning each of the refined models, by incorporating segmentation clustering, treating this as a multitask learning problem. Extensions to image and text data could also be considered, possibly utilizing pre-trained models in place of the base models in Algorithm 1. These problems however would involve further consideration of the trade-off between domain invariance properties and refinement for each segment. Lastly, our framework of studying a segmented distribution shift may have further connections with algorithmic fairness, as we avoid the problem of large segments dominating the training of the model. However, recent work explores the impact of distribution shift on fairness metrics and may be worth further exploration in segmented settings (Jiang et al., 2024).
Acknowledgements
The authors thank Sai Sumanth Miryala, Tao Wu, Xiaolu Xiong, and Andy Yu from the People Data Platform team for their support and feedback throughout this research.
@@ A.1 Methods for Learning Importance Weights
We can apply standard techniques for learning the importance weights of the function class. Firstly, we consider learning covariate shift weights through discriminative learning, i.e. learning a probabilistic classifier that discerns between the weights (Bickel et al., 2009). Let $T=\{0,1\}$ denote whether a sample is within the training or test distribution letting $P(\cdot)=\mathbb{P}(\cdot|T=0)$ and $Q(\cdot)=\mathbb{P}(\cdot|T=1)$, then the importance weights can be expressed up to proportion as
$w(x)=\frac{dQ(x)}{dP(x)}\propto\frac{\mathbb{P}(T=1|x)}{\mathbb{P}(T=0|x)}.$
@@
Secondly, we consider Penalized Risk Minimization (Nguyen et al., 2010). This method involves the estimation as follows:
$\widehat{w}_{s}(x)$ $=\widehat{g}_{s}(x)$
$\widehat{g}_{s}$ $=\text{argmin}_{g\in\mathcal{G}}\frac{1}{n_{s,tr}}\sum_{i\in N_{s}1}g(X_{i})$
$-\frac{1}{n_{s,test}}\sum_{j\in N^{\prime}_{s}}^{n}\log(g(X^{\prime}_{j}))+\frac{\gamma_{n}}{2}\widetilde{I}^{2}(g)$
where $\mathcal{G}$ is a reproducing Kernel Hilbert space, $N_{s}$ and $N^{\prime}_{s}$ refer to the indices corresponding to training and test indices in segment $s$ and $\widetilde{I}$ is a non-negative measure of complexity for $g$ such that $\widetilde{I}(w)<\infty$ (for example, $\widetilde{I}(g)=\lVert g\rVert_{\mathcal{F}}$).
Another common approach to estimating importance weights is via Kernel Mean Matching (KMM) (Gretton et al., 2012). Given a positive definite kernel function $k(\cdot,\cdot)\to\mathbb{R}$ let $G_{ij}=k(X_{i},X_{j})$ and let $\kappa_{i}=\frac{n_{tr}}{n_{te}}\sum_{j=1}^{n_{te}}k(X_{i},X_{j}^{\prime})$. Then kernel mean matching computes a set of weights $\{w_{i}\}_{i=1}^{n_{tr}}$
@@
Lastly, when correcting for label shift one can use the Black Box Shift Estimation (BBSE) of (Lipton et al., 2018). This involves using a predictor $h(\cdot)$, and denote $\widehat{Y}=h(X)$. The corresponding importance weights $\mathbf{\tilde{w}}\in\mathbb{R}^{K}$ where $\tilde{w}(y)=\frac{Q(y)}{P(y)}$, can be defined through a linear equation of the expected confusion matrix under the training distribution $\mathbf{C}$, where $\mathbf{C}_{ij}=P(\widehat{Y}=i,Y=j)$ and the predicted class frequencies on the test distribution $\mathbf{\mu}_{Q,h}\in\mathbb{R}^{K}$ where $\mu_{Q,h,j}=Q(\widehat{Y}=j)$. We differentiate the importance weights for each class $\mathbf{\tilde{w}}\in\mathbb{R}^{K}$ from $\mathbf{w}\in\mathbb{R}^{n}$, the weight vector for each training observation using the tilde annotation. The importance weights can then be defined by the solution to the following:
@@
Using the training labels, one can estimate $\mathbf{\widehat{\mu}}_{Q,h}$ and $\mathbf{\widehat{C}}$ using empirical means. As long as $\mathbf{C},\mathbf{\widehat{C}}$ are invertible, then one can estimate the importance weights of the class
@@
In order to prove Theorem 5.3 we first introduce a series of supporting lemmas, after which we proceed with the proof of the theorem in Appendix B.3.
B.1 Supporting Lemmas
Lemma B.1 (Best Linear Combination As A Prior).
Let $\{h_{m}\}_{m=1}^{M}$ be a collection of base models contained in some function space $h_{m}\in\mathcal{F}$. Let $s$ denote the target segment of interest, and let $N_{s}\subset\{1,2,n_{tr}\}$ be the set of indices of the training data corresponding to segment $s$ with $|N_{s}|=n_{s}$. Let $\beta\in\mathcal{B}\subset\mathbb{R}^{M}$ denote the parameters for a linear combination of $h_{m}$. Let $\mathcal{B}$ be a convex set in the parameter space. Let $\beta^{*(s)}$ denote the optimal linear combination satisfying $\beta^{*(s)}=\text{argmin}_{\beta\in\mathcal{B}}R_{P}[\mathbf{h}^{\intercal}\beta|s]$ and the empirical estimate $\widehat{\beta}^{(s)}=\text{argmin}_{\beta\in\mathcal{B}}R[\mathbf{h}^{\intercal}\beta|\mathbf{Y},\mathbf{X},s]$. Denote the following random variables and parameters derived from the loss function and the base predictors:
For a universal constant $C$
Lemma B.1 follows from a simple application of the results from (Ostrovskii & Bach, 2021). The main idea we exploit is that learning a simple linear combination is easier than learning in the larger function class $\mathcal{F}$. Since we assume that the total number of base estimators is not increasing, then $M_{eff}$, $M_{F},M$ and $\tilde{\lambda}_{1}$ are all constants and we recover the stage 1 model at a rate of $n_{s}^{-1/2}$. This therefore, allows us to learn this prior faster than that of the stage 2 model.
B.2 Proof of Lemma B.1
Proof.
Recall $h_{m}:\mathcal{X}\to\mathcal{Y}$ are the set of base-models, and let $\mathbf{h}$ denote the vector of such functions. Denote a linear combination of these with weights $\beta$ such that $\mathbf{h}^{\intercal}\beta=\sum_{m=1}^{M}h_{m}\beta_{m}$.
The first step is an immediate application of Theorem 1.1 of (Ostrovskii & Bach, 2021). This theorem is applied as follows. Let $\rho$ denote the minimum value such that $\Sigma\preceq\rho H$. Then for all $n_{s}\geq\rho M_{eff}M\log(eM/\delta)$
$R[\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}|s]-R[\mathbf{h}^{\intercal}\beta^{*(s)}|s]\lesssim(\widehat{\beta}^{(s)}-\beta^{*(s)})^{\intercal}H^{2}(\widehat{\beta}^{(s)}-\beta^{*(s)})\lesssim\frac{M_{eff}\log(e/\delta)}{n_{s}}$
where $\lesssim$ denotes less than or equal to, up to a universal constant $C$.
Further let $\tilde{\lambda}_{1}$ denote the smallest eigenvalue of $H$. Then with probability at least $1-\delta$
$\lVert\widehat{\beta}^{(s)}-\beta^{*(s)}\rVert_{2}^{2}\lesssim\frac{M_{eff}\log(e/\delta)}{\lambda^{2}_{1}n_{s}}.$
$\lVert\mathbf{h}^{\intercal}\beta^{*(s)}-\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}\rVert_{\mathcal{F}}$ $\leq\sum_{m=1}^{M}\lVert h_{m}(\beta_{m}^{*(s)}-\widehat{\beta}_{r})\rVert_{\mathcal{F}}$
We next consider the following lemma, first introduced in (Reddi et al., 2015) with its few line proof for completeness.
Lemma B.2 (Lemma 2 of (Reddi et al., 2015)).
$\sup_{f:\lVert f-f_{0}\rVert\leq\nu}\left|R[f|\mathbf{Y},\mathbf{X},s;w_{s}]-R_{Q}[f|s]\right|$ $\leq\eta L\left(\frac{2\nu}{n_{s}}\sqrt{\text{tr}(\mathbf{K}_{s})}+3\sqrt{\frac{\log(2/\delta)}{n_{s}}}+\frac{1}{n_{s}}\sqrt{\sum_{i\in N_{s}}f_{0}^{2}(X_{i})}\right)$
Proof.
[…]
where $\sigma_{i}$ are independently drawn uniform {±$\{\pm 1\}$ random variables.
[…]
Lastly, we include an upper bound on the weights.
Lemma B.3.
Let $\widehat{w}_{s}$ be an estimator of the true importance weights. Then
$\sup_{f\in\mathcal{F}}|R[f|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]-R[f|\mathbf{Y},\mathbf{X},s,w_{s}]|\leq L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
Proof.
The proof is a straightforward application of Holder’s inequality.
$\sup_{f\in\mathcal{F}}|R[f|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]-R[f|\mathbf{Y},\mathbf{X},s,w_{s}]|$ $=\sup_{f\in\mathcal{F}}|\frac{1}{n_{s}}\sum_{i\in N_{s}}\ell(Y_{i},f(X_{i}))(\widehat{w}_{s}(X_{i},Y_{i})-w_{s}(X_{i},Y_{i}))|$
$\leq\sup_{f\in\mathcal{F}}\text{max}_{i\in N_{s}}|\ell(Y_{i},f(X_{i}))|\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
$\leq L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
B.3 Proof of Theorem 5.3
@@ B.1 Supporting Lemmas
Recall the two-stage multiply robust estimator
$\widehat{f}_{MR}^{(s)}$ $=\text{argmin}_{f\in\mathcal{F}}R[f|\mathbf{Y},\mathbf{X},s;\widehat{w}_{s}]+\lambda^{\prime}\lVert\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}-f\rVert_{\mathcal{F}}$
$=\text{argmin}_{f:\lVert\mathbf{h}^{\intercal}\widehat{\beta}^{(s)}-f\rVert_{\mathcal{F}}\leq\nu^{\prime}}R[f|\mathbf{Y},\mathbf{X},s;\widehat{w}_{s}]$
where the constraint value $\nu^{\prime}$ denotes the corresponding value to the regularization parameter $\lambda^{\prime}$.
@@
$\sup_{f:\lVert f-\mathbf{h}^{\intercal}\widehat{\beta}\rVert\leq\nu^{\prime}}\left|R[f|\mathbf{Y},\mathbf{X},s;w_{s}]-R_{Q}[f|s]\right|$ $\leq B(n_{s},\delta)$
$B(n_{s},\delta)$ $:=\eta L\left(\frac{2\nu^{\prime}}{n_{s}}\sqrt{\text{tr}(\mathbf{K}_{s})}+3\sqrt{\frac{\log(2/\delta)}{n_{s}}}+\frac{1}{n_{s}}\sqrt{\sum_{i\in N_{s}}(\mathbf{h}^{\intercal}\widehat{\beta})^{2}(X_{i})}\right)$
@@
$|R_{Q}[f|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]-R_{Q}[f|\mathbf{Y},\mathbf{X},s,w_{s}]|\leq C(n_{s},\widehat{w}_{s}):=L\frac{\lVert\widehat{w}_{s}-w_{s}\rVert_{1}}{n_{s}}$
@@
$R_{Q}[\widehat{f}_{MR}^{(s)}|s]$ $\leq R_{Q}[\widehat{f}_{MR}^{(s)}|\mathbf{Y},\mathbf{X},s,w_{s}]+B(n_{s},\delta)$
$\leq R[\widehat{f}_{MR}^{(s)}|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]+B(n_{s},\delta)+C(n_{s},\widehat{w}_{s})$
$\leq R[f^{*(s)}_{\nu_{q,s}}|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]+B(n_{s},\delta)+C(n_{s},\widehat{w}_{s})$
$\leq R[f^{*(s)}_{\nu_{q,s}}|\mathbf{Y},\mathbf{X},s,w_{s}]+B(n_{s},\delta)+2C(n_{s},\widehat{w}_{s})$
@@
The third inequality follows from the fact $R[\widehat{f}_{MR}^{(s)}|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]\leq R[f^{*(s)}_{\nu_{q,s}}|\mathbf{Y},\mathbf{X},s,\widehat{w}_{s}]$
@@
B.4 Bounds on estimating the importance weights
Our upper bound in Theorem 5.3 relies on estimating the importance weights $\lVert\widehat{w}_{s}-w_{s}\rVert_{1}$. We focus on two upper bounds, one for covariate shift and one for label shift.